What is the value of this trigonometry  theta
 
 
Sin37°
Sin53°
Cos37°
Sin53°

Grade:11

1 Answers

Snehal Jadhav
5 years ago
Overhere, these are standard results:
sin(37)=3/5
 
sin(53)=4/5
 
cos(37)=4/5
 
cos(53)=3/5
 
tan(37)=3/4
 
tan(53)=4/3